# Entangled in Representations: Mechanistic Investigation of Cultural Biases in Large Language Models

### v1

@@ Appendix A Dataset Details
In our experiments, we utilize BLEnD Myung et al. (2024) and CAMeL-2 Naous and Xu (2025). We provide their brief data statistics and characteristics in Table 4.
A.1 Country Groups
As mentioned in Section 3.2, our work conducts experiments that focus on 14 countries classified in two dimensions. We compare three groups based on the level of language resource Joshi et al. (2020) and six groups based on the continental region. Table 3 shows the country entities that correspond to each group.
@@ Appendix B Example of Prompts
Appendix C CF Score Results
@@
In Figure 6, we show the CF score results with Llama-3.1, aya-expanse, Qwen2.5 on BLEnD and CAMeL-2. As mentioned in §6.2, we exclude the countries with CF scores lower than the average CF score across all countries.
(a) BLEnD, aya-expanse
(b) BLEnD, Qwen2.5
(c) CAMeL-2, aya-expanse
(d) CAMeL-2, Qwen2.5
(e) CAMeL-2, Llama-3.1
@@
Appendix D Attention Contributions
In cases where an option consists of multiple tokens, we follow the approach of Yuksekgonul et al. (2024), taking the maximum attention contribution score among the component tokens. When the attention scores are averaged by samples to examine its general patterns, simply averaging can be sensitive to extreme values or samples in which attention scores are globally biased toward higher or lower magnitudes. This variability can potentially reduce the generalizability of the results. To address this, we applied normalization (z-score normalization) per sample to the attention contribution scores, such that the scores within each sample have a mean of 0 and a standard deviation of 1.

### v2

@@ A.1 Culture Groups
BLEnD-Resource BLEnD-Region
Hypothesis 1 Hypothesis 2 Hypothesis 3 Hypothesis 1 Hypothesis 2 Hypothesis 3
Llama-3.1 low mid avg. 0.008 low low mid avg. 0.0049 low
avg. 0.0034 0.6328 mid 0.2256 avg. 0.1066 avg. 0.0004 0.931 mid 0.0016 avg. 0.7306
mid 0.0107 - low 0.076 mid 0.1008 mid 0.0001 - low 0.8067 mid 0.8815
aya-expanse low mid avg. 0.009 low low mid avg. 0.0124 low
avg. 0.0 0.936 mid 0.1477 avg. 0.0946 avg. 0.0007 0.8161 mid 0.0086 avg. 0.4606
mid 0.0 - low 0.2307 mid 0.0548 mid 0.0007 - low 0.5762 mid 0.6784
Qwen2.5 low mid avg. 0.0 low low mid avg. 0.0 low
avg. 0.026 0.461 mid 0.22 avg. 0.3283 avg. 0.0 0.9722 mid 0.0122 avg. 0.1162
mid 0.0796 - low 0.0215 mid 0.1636 mid 0.0 - low 0.0307 mid 0.1076
Table 7: P-values from two-proportions Z-tests with three different hypotheses. Bolded texts mean that the test supports the hypothesis.
A.2 Cultural MCQ with hard negatives
The creation of BLEnD-Resource and BLEnD-Region involves random selection among possible options (Sec. 3.2.3). Table 4 presents descriptive statistics including mean, standard deviation, and 95% confidence interval of the number of selected culture groups for each option type. This demonstrates that options are sampled to achieve a balanced representation of cultural groups. A gold answer option type refers to the culture group from which the option originates. An option type of hard negative refers to a culture group from which an option selected as a hard negative originates. High, mid, and low denote the number of selected countries from three resource groups.
@@ Appendix B Example of Prompts
Appendix C Details for Culturescope
BLEnD CAMeL-2
Llama-3.1 English 0.325 $\pm$ 0.01 0.239 $\pm$ 0.07
Spanish 0.277 $\pm$ 0.07 -
Irrelevant 0.149 $\pm$ 0.09 0.155 $\pm$ 0.08
aya-expanse English 0.349 $\pm$ 0.1 0.242 $\pm$ 0.07
Irrelevant 0.127 $\pm$ 0.07 0.141 $\pm$ 0.06
Qwen-2.5 English 0.344 $\pm$ 0.11 0.227 $\pm$ 0.07
Arabic - 0.173 $\pm$ 0.07
Irrelevant 0.141 $\pm$ 0.09 0.148 $\pm$ 0.07
Table 8: Average and standard deviations of the similarity scores from the filtering stage of the Culturescope. We round the average off to two decimal places.
CAMeL-2
Relevance Evaluation
Llama-3.1 aya-expanse Qwen2.5
Baseline 0.7126 0.7019 0.6675
Cultural Prompting 0.6799 0.7282 0.7148
CANDLE 0.5640 0.6955 0.6473
CultureScope 0.6519 0.7169 0.6659
Table 9: We present the relevance evaluation of Culturescope on CAMeL-2 in English. We highlight the best-performing method in bold and the second-best with an underline.
low vs high low vs mid
Llama-3.1 1.11e-09 0.159
aya-expense 8.55e-10 9.69e-13
Qwen2.5 1.75e-07 5.54e-03
Table 10: p-value of Student’s t-test for each model on low-resource target samples from BLEnD-Resource, based on attention scores. Comparisons were conducted between low and mid, low and high, all from $C_{resource}$, which are all hard negative options.
During the filtering step in the Culturescope, we use three different semantic search models for each language. For English, we use a DeBERTa based model, trained for semantic search [^1] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. For Arabic, we use the bert-base model finetuned for semantic search in Arabic [^2] https://huggingface.co/akhooli/Arabic-SBERT-100K. Lastly, for Spanish, we use the bert-base model finetuned for semantic search in Spanish [^3] https://huggingface.co/hiiamsid/sentence_similarity_spanish_es.
In Table 8, we show the threshold values selected for each dataset and language for the filtering stage. The values are selected by the average of cosine similarity score across all the generated concepts.
Figure 6 describes the average number of generated knowledge per layer. This shows that Culturescope successfully generates relevant cultural knowledge from layers, especially from early layers.
Appendix D Statistical Significance Tests
D.1 Testing on MCQ Results
To evaluate the statistical significance of our findings in Sec. 6.2.1, we set three alternative hypotheses and test with a two-proportions Z-test  [^4] https://www.statsmodels.org/dev/generated/statsmodels.stats.proportion.proportions_ztest.html. Table 7 presents p-values from the list of Z-tests. avg., mid, and low follow the definition from Sec. 6.2.1. The Z-test supports the alternative hypothesis when the p-value is lower than 0.05. Hypothesis 1 is that the accuracy from the avg. and the mid is higher than the accuracy from the low. The p-values in all models on both datasets support this hypothesis. Hypothesis 2 is that in each level, when an LLM predicts incorrectly, it prefers the biased option over the other options. This hypothesis is supported in avg. and mid groups, which highlights our finding on less susceptibility of low-resource cultures with cultural biases. Lastly, hypothesis 3 is that the model predicts biased options more in the low group than in the avg. and mid groups. This hypothesis is rejected in all tests, confirming the statistical significance of lower cultural bias in the low group.
Llama-3.1 aya-expanse Qwen2.5
Target Target vs EUR Target vs AME Target vs EUR Target vs AME Target vs EUR Target vs AME
S-AS 0.0570 0.199 0.614 0.596 0.00109 0.0413
E-AS 0.864 0.403 0.0851 0.787 0.799 0.609
W-AS 0.827 0.476 0.787 0.829 0.329 0.475
EUR 4.22e-03 0.216 2.26e-06 1.89e-04 0.160 0.104
Table 11: p-value of Student’s t-test for each model on non-Western targeted samples from BLEnD-Region, based on attention scores. Comparisons were made between each sample’s target region and America, and between each sample’s target region and Europe, all within $C_{region}$, which represents hard negative options.
D.2 Testing on the Results for Attention Contribution Score
As discussed in Sec.  6.1.2, our analyses suggest that high-resource bias and Western-dominance bias are strongly internalized within LLM representations, as visualized in the attention heatmaps in Figure 4 and Figure 8. To further validate these observations, we conduct Student’s t-tests [^5] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.ttest_ind.html on the attention scores of BLEnD-Resource and BLEnD-Region.
To examine bias patterns, we conduct p-value analyses on attention scores using Student’s t-tests. For high-resource bias, low-resource targets ($c_{low}$) are compared against mid- and high-resource ($c_{mid}$, $c_{high}$) samples within $C_{resource}$, with most p-values less than 0.05, indicating statistically significant differences (Table 10). For Western-dominance bias, for non-Western regions (South Asia, East Asia, West Asia) targeted samples, non-Western regions ($c_{sameregion}$) are compared against Western regions (Europe, America) ($c_{region1}$, $c_{region2}$) within $C_{region}$. While fewer comparisons yield significance, several regions with large mean differences show p-values less than 0.05 (Table 11), suggesting Western-dominance bias is present but less uniformly across regions.
In summary, the results confirm that high-resource bias is consistently strong in BLEnD-Resource, while Western-dominance bias is also evident in BLEnD-Region, though less uniformly across all regions.
Appendix E CF Score Results
@@
In Figure 7, we show the CF score results with Llama-3.1, aya-expanse, Qwen2.5 on BLEnD and CAMeL-2. As mentioned in Sec. 6.1.1, we exclude the countries with CF scores lower than the average CF score across all countries.
1. Inspection prompts
Generate associated words, Syria, Oman, Jordan, Qatar, West Asia, Turkey, Israel, Lebanon, …, Leonardo DiCaprio, Tom Cruise, Kate Winslet, Brad Pitt, Actor, …, Samsung, Cell Phone, TV, Apple, Nokia, South Korea, Electronics, …, x
2.Open-ended QA Prompts
2-1. Baseline
BLEnD
[…]
BLEnD
[…]
BLEnD
[…]
Figure 5: Prompt templates used for open-ended QA evaluations.
(a) BLEnD, English
(b) CAMeL-2, English
Figure 6: Average number of generated knowledge per layer.
(a) BLEnD, English, aya-expanse
(b) BLEnD, English, Qwen2.5
(c) BLEnD, Spanish, Llama-3.1
(d) CAMeL-2, English, Llama-3.1
(e) CAMeL-2, English, aya-expanse
(f) CAMeL-2, English, Qwen2.5
(g) CAMeL-2, Arabic, Qwen2.5
@@ 2-1. Baseline
Appendix F Attention Contributions
In cases where an option consists of multiple tokens, we follow the approach of Yuksekgonul et al. (2024) and take the maximum attention contribution score across the component tokens. When the attention scores are averaged by samples to examine their general patterns, simply averaging can be sensitive to extreme values or samples in which attention scores are globally biased toward higher or lower magnitudes. This variability can potentially reduce the generalizability of the results. To address this, we applied normalization (z-score normalization) per sample to the attention contribution scores, such that the scores within each sample have a mean of 0 and a standard deviation of 1.